OpenStudy (anonymous):

4/5= which ratio will form a proportion with 4/5 a.8.8/5.5 b.88/100 c.6/7.5 d.18/20

OpenStudy (anonymous):

c

OpenStudy (anonymous):

since 6 = 4 * 1.5 and 7.5 = 5 * 1.5